OpenStudy (anonymous):

Simplify 32^3/5 –7 20 19.2 8

OpenStudy (anonymous):

its actually (2^5)3/5 =8

OpenStudy (anonymous):

\[\sqrt[5]{32^3}\]

OpenStudy (anonymous):

(32)^3/5 = (2^5) ^3/5 = 2^3 = 8